If a 23.7 mL solution of H2SO4 has a concentration of 0.100 M, what is its pH? Also, how many grams of H2SO4 are in the solution?
1
Expert's answer
2020-05-04T01:53:20-0400

Solution:

Sulfuric acid - H2SO4 - strong diprotic acid, completely dissociates into ions in solution.

The dissociation equation is written as:

H2SO4 = 2H+ + SO42-

According to the equation: CM(H2SO4) = [H+]/2.

[H+] = 2 × CM(H2SO4) = 2 × (0.100 M) = 0.200 M


pH = -log[H+] = -log(0.200) = 0.69897 = 0.70

pH = 0.70


Molarity of H2SO4 = Moles of H2SO4 / Volume of solution

Moles of H2SO4 = n(H2SO4) = CM(H2SO4) × V(solution)

n(H2SO4) = (0.100 M) × (0.0237 L) = 0.00237 mol


Moles of H2SO4 = Mass of H2SO4 / Molar mass of H2SO4

The molar mass of H2SO4 is 98.079 g/mol.

Mass of H2SO4 = m(H2SO4) = n(H2SO4) × M(H2SO4)

m(H2SO4) = (0.00237 mol) × (98.079 g/mol) = 0.2324 g

Mass of H2SO4 is 0.2324 g


Answer:

pH = 0.70

0.2324 grams of H2SO4 are in the solution.